# Branch length statistics in phylogenetic trees under constant-rate birth-death models

### v2

@@ Abstract
Phylogenetic trees represent the evolutionary relationships between extant lineages, where extinct or non-sampled lineages are omitted. Extending the work of Stadler and collaborators, this paper focuses on the branch lengths in phylogenetic trees arising under a constant-rate birth-death model. We derive branch length distributions of phylogenetic branches with and without random sampling of individuals of the extant population under two distinct statistical scenarios: a fixed age of the birth-death process and a fixed number of individuals at the time of observation. We find that branches connected to the tree leaves (pendant branches) and branches in the interior of the tree behave very differently under sampling: pendant branches get longer as the sampling probability is decreased, whereas the distribution of interior branches quickly reaches an asymptotic regime.
@@ I Introduction
The reconstructed evolutionary process is central to understanding evolutionary relationships and dynamics. This reconstruction is based on extant genetic data and yields a phylogenetic tree, which represents the evolutionary relationships among a set of species, genes, or cells. Crucially, all lineages that eventually went extinct (or were not sampled) are missing from such a tree. This affects the statistics of the reconstructed tree: Under a stochastic birth-death process, species go extinct at a specific rate (or cells die, in a cellular system). Looking at the reconstructed tree, however, only speciation events are seen. To compare empirical phylogenetic trees with models of evolution, it is crucial to understand the link between an evolutionary process and the reconstructed process.
@@
The reconstructed evolutionary process has been studied extensively, with early work of Nee, May, and Harvey Nee et al. (1997) giving the probability density in time for each speciation event in a reconstructed tree from a birth-death process with constant rates. Such a birth-death model describes a population of species, individuals, or cells, which divide independently at a fixed rate $\lambda$ and die independently at a fixed rate $\mu$. (For species, birth events describe the emergence of a new species, whereas death events correspond to extinctions.) A constant-rate birth-death model serves as a simple model of neutral evolution.
@@
Stadler and collaborators have extended the work by Nee, May, and Harvey Nee et al. (1997) to trees reconstructed from individuals that were sampled from a population Stadler (2009) and to populations conditioned on a particular size Gernhard (2008). The resulting statistics of speciation events have been used to infer trees and their characteristics Stadler (2013b); MacPherson et al. (2022), and is the cornerstone of widely-used computational packages such as BEAST Bouckaert et al. (2019) and TreeTime Sagulenko et al. (2018). Many extensions of the basic constant-rate birth-death process have been developed, for instance, to account for rates that are variable through time or vary across the tree (selection) Stadler (2013b).
@@
In Stadler and Steel (2012); Mooers et al. (2012) Stadler, Steel and collaborators focus on a particular aspect of tree statistics, namely the distribution of branch lengths (in calendar time). They calculate the probability density function for the length of the so-called pendant branches (the lowest branches leading to the leaves of the tree) for trees reconstructed from a constant-rate birth-death process. For the particular case of a Yule process (a birth-death process with the rate of death set to zero), they also give the probability density function for the length of interior branches.
@@
In this paper, we revisit the problem of the distribution of branch lengths with a particular focus on the lengths of interior branches and the distribution of branch lengths under sampling of extant species. Two statistical scenarios need to be distinguished: In the first scenario (scenario i)), the time interval between the start of the birth-death process, starting with a single individual, and the time of observation (the end of the birth-death process) is known, but the number of extant species at the end of that process is a fluctuating random variable. In the second scenario (scenario ii)), one conditions on the number of individuals at the end of the process, but the time interval between the start and end of the process is unknown.
@@
In Section II, we recapitulate the relevant properties of a constant-rate birth-death process and in Sections III-V, we discuss a simple but effective approach to determine branch length distributions under the first scenario used by Paradis Paradis (2016). We write down concrete expressions for branch length distributions in the limit of a long-running birth-death process, focussing on pendant branches in Section III, on branches one and two levels above a pendant branch in Section IV, and on generic branches picked randomly in the interior of a phylogenetic tree in Section V. Section VI briefly discusses the statistics of the number of generations along branches. In Section VII we generalize these results to a scenario where only a finite fraction of the extant population is sampled. We find that pendant and interior branch lengths behave very differently under sampling: while pendant branches become longer and grow without limit as the sampling probability approaches zero, the distribution of interior branch lengths quickly reaches an asymptotic limit as the sampling probability decreases. In Section VIII we turn to the second scenario (conditioning on the number of extant species at the time of observation) and introduce a novel approach based on tracking the birth-death process first backwards in time from the time of observation to the emergence of a particular branch and then again forwards to the time of observation. For pendant branches, we recover the results of Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and generalize these results to interior branches.
@@ II Dynamics of a birth-death process
We consider a birth-death process as introduced by Feller Feller (1939): individuals in a population duplicate with a constant rate of birth $\lambda$ and die with a constant rate of death $\mu$. The dynamics of the population size are described by a Markov process, with the probability $p_{n}(t)$ of the population consisting of $n$ individuals at time $t$ following a master equation
@@
It turns out to be convenient for the analysis of branch lengths to consider times relative to the time of observation $T$. We denote the length of the interval between the intermediate time $t$ and the observation time $T$ by $\tau_{e}=T-t$. We further define $\tau=T-\tau_{e}$ as the length of the interval between $\tau_{e}$ and the time of the clade’s origin. Specifically, we denote as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{(T)}_{1}(t=T-\tau_{e})$ the probability that a clade started at time $\tau_{e}+\tau$ in the past, and contains at time $\tau_{e}$ in the past only a single individual that could have extant descendants at the time of observation. Again, setting $\tau_{e}=0$ results in $p_{1}^{(\tau_{e}=0)}(\tau)\equiv p_{1}(\tau)$. For the case of $\tau_{e}>0$ the integrating factor becomes $-\int_{0}^{T-\tau_{e}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]=-\int_{0}^{\tau}\operatorname{\mathrm{d}\!}\tau^{\prime}\left[\mu+\lambda\left(1-2p_{0}(\tau_{e}+\tau^{\prime})\right)\right]=\log\left(\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\right)$ and we obtain
@@ III Branch length distribution: pendant branches
The results from section II can be used directly to determine the statistics of branch lengths of a phylogenetic tree. In this section, we apply them to the so-called pendant branches, which end in a leaf of the phylogenetic tree. We use the approach of Paradis Paradis (2016) to determine branch length distributions and write down concrete expressions for the probability densities. This section and the following are concerned with complete sampling (where the phylogenetic tree is constructed from the entire population alive at a particular time since the start of the birth-death process), in Section VII we will turn to the case where only a finite fraction of the population is available to reconstruct a tree from.
@@
For a finite time since the start of the birth-death process, $\tau$ has a finite support and hence $P_{\text{pend}}(\tau)$ cuts off to zero when $\tau$ is larger than the time $T$ since the start of the birth-death process. The normalizing factor of (10) is then found to be $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(\lambda(\lambda-(\lambda-2\mu)\operatorname{\mathrm{e}}^{-2(\lambda-\mu)T}-2\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))$. We focus on the particularly simple case of an asymptotically long running time $T\to\infty$. In this case, the cutoff moves to infinity and we find the normalizing factor in (10) to be one, making (10) a proper probability density function. We will focus on the asymptotic case $T\to\infty$ also when treating branches in the tree’s interior.
@@ IV Branch length distribution: interior branches above pendant branches
Equation (11) gives the joint probability density function of branch lengths of pendant and level-two branches up to a normalization factor, which again turns out to be one. Figure 2D compares this result to numerical simulations. A population with birth rate $\lambda=1$ and death rate $\mu=0.75$ was grown from a single individual until time $T=25$ (where the population contained 3567 individuals), and the phylogenetic tree was reconstructed. The lengths of level-two branches were collected in a histogram and compared to the joint probability $P_{\text{level}}(\tau_{1},\tau_{2})$ given by (11) marginalized with respect to $\tau_{1}$.
@@ V Branch length distribution: interior branches
So far, we have again only considered the asymptotic case of $T\rightarrow\infty$. For finite $T$, the probability of $\tau_{s}>T$ is 00, resulting in a normalization factor of $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(a(1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))^{2}$ for (13).
VI Number of generations along branches
Birth events where one of the two newly founded clades dies out do not lead to bifurcations in the reconstructed tree. Such birth events are not visible from the reconstructed tree alone, yet they contribute to the number of generations along a lineage. In this section, we discuss the statistics of the number of generations along a particular branch of a phylogenetic tree.
The mean number $i$ of generations along a particular branch starting at time $\tau_{s}$ and ending at time $\tau_{e}$ (not counting the founding event of the branch) has been calculated by Bokma, van den Brink, and Stadler Bokma et al. (2012) as
$\langle i\rangle(\tau_{s},\tau_{e})=2\lambda\int_{\tau_{e}}^{\tau_{s}}\operatorname{\mathrm{d}\!}\tau\,p_{0}(\tau)=2\left[\mu\tau-\log\left(\lambda-\mu e^{-(\lambda-\mu)\tau}\right)\right]_{\tau_{e}}^{\tau_{s}}$ (15)
based on birth events occurring along a branch at rate $\lambda$ and the probability $p_{0}(\tau)$ that one given additional lineage has died out by time $\tau$ given by (2). The factor $2$ arises because either of the two offspring can be the one to die out. The number of generations is Poisson distributed (and hence fully characterized by the mean (15)), because successive birth events are statistically independent Bokma et al. (2012).
Together with the densities of branch lengths from the previous sections, we can now compute the probability distribution for the number of generations along a randomly chosen branch. In the case of pendant branches and asymptotically old trees, it is given by
$P_{\text{pend}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau\,P_{\text{pend}}(\tau)\operatorname{\mathrm{e}}^{-\langle i\rangle(\tau,0)}\frac{\left[\langle i\rangle(\tau,0)\right]^{i}}{i!}\ .$ (16)
Similarly, the probability of finding $i$ generations along a randomly chosen interior branch is
$P_{\text{int}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau_{s}\int_{0}^{\tau_{s}}\operatorname{\mathrm{d}\!}\tau_{e}\,P_{\text{int}}(\tau_{s},\tau_{e})\operatorname{\mathrm{e}}^{-\langle i\rangle(\tau_{s},\tau_{e})}\frac{\left[\langle i\rangle(\tau_{s},\tau_{e})\right]^{i}}{i!}\ .$ (17)
VII Finite sampling
@@
$P_{\text{int}}^{(\rho)}(\tau)$ $=(\lambda-\mu)^{2}\frac{2\lambda\rho-\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left(\mu-\lambda(1-\rho)\right)}{\left(\lambda(1-\rho)-\mu+\lambda\rho\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}\right)^{2}}$ (23)
@@
ARefer to caption BRefer to caption                   C Refer to caption
Figure 3: Pendant and interior branches at different sampling probabilities $\rho$. (A) We compare the analytical results for the probability density of the length of pendant and interior branches under sampling ((21) and (23) respectively, solid blue lines) to numerical simulations. For the latter, we show the distribution of branches from 100 trees. For every tree, we simulated the growth of a population with $\lambda=1$ and $\mu=0.75$ over time $T=40$, restarting when the population died out or the final population size was not within $2\times 10^{4}$ and $3\times 10^{5}$ (grey bars). The first row of plots shows the branch length distribution of the full trees. For the next row of plots, individuals were sampled with probability $\rho=0.1$ and the corresponding trees were reconstructed. For the third row, each of the previous samples was sampled again with probability $0.1$, corresponding to an overall sampling probability of $\rho=0.01$. The same was repeated for the forth row of plots, yielding $\rho=0.001$. Every tree contributes equally to the empirical distributions depicted. The first column shows the distributions of pendant branch lengths, which shift to longer branches as $\rho$ decreases. The second column of plots shows the distribution of interior branches, quickly reaching the asymptotic distribution (24) (red dashed line) as $\rho$ becomes small. (B) The distribution of the number of generations along interior branches in the asymptotic case of small sampling probabilities (25). Different colours represent different relative death rates $\mu/\lambda$. (C) A toy example of how sampling affects the length of pendant branches, see text.
@@
The difference between pendant and interior branches under sampling has a simple qualitative explanation, which we explore within a toy problem. Consider a reconstructed tree which is perfectly balanced; both clades defined by each bifurcation have the same size. An example is shown in Fig. 3C. Then a fraction $\rho$ of the extant nodes is sampled. Under sampling, a pair of pendant branches linked to a shared ancestor is defined by the birth of a clade (shown in blue in Fig. 3C) whose size $n$ at the time of observation is defined by $n\rho\approx 2$ (so two of its members are expected to be sampled). The number of branch levels between this founding birth event and the time of observation is $\log_{2}(n)$, so the pendant branch lengths are increased under sampling by a factor of $1-\log_{2}\rho\approx-\log_{2}\rho$ relative to the fully sampled case. (This argument assumes that on the fully sampled reconstructed tree, pendant and interior branches are approximately of the same length.) On the other hand, branches that are interior in the reconstructed tree (i.e. lie above birth event shown in blue in Fig. 3C) have their lengths unchanged by sampling. The distribution of interior branches only changes because under sampling, fewer and older branches are counted as interior. Older branches tend to be a little longer (see equation (13)), because any additional lineages that arose along them had more time to die out until the time of observation.
Following Section VI and equations (21)-(23), we also compute the distribution of generations along branches of a sampled tree. Crucially, since the mean number of generations (equation (15)) along a branch of length $\tau$ converges to $2\mu\tau$ as the branch moves further away from the time of observation (as interior branches do when taking the limit $\rho\rightarrow 0$), we obtain the probability of finding $i$ generations along an interior branch in this limit
@@
$P^{(\rho\rightarrow 0)}_{\text{int}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau\,P^{(\rho\rightarrow 0)}_{\text{int}}(\tau)\operatorname{\mathrm{e}}^{-2\mu\tau}\frac{(2\mu\tau)^{i}}{i!}=(2\mu)^{i}(\lambda-\mu)(\lambda+\mu)^{-i-1}\ .$ (25)
This asymptotic result is depicted for different relative death rates $\mu/\lambda$ in Figure 3B.
VIII Branch length distributions conditioned on the population size at the time of observation

### v3

# Branch length statistics in phylogenetic trees under constant-rate birth-death dynamics

@@ Abstract
Phylogenetic trees represent the evolutionary relationships between extant lineages, where extinct or non-sampled lineages are omitted. Extending the work of Stadler and collaborators, this paper focuses on the branch lengths in phylogenetic trees arising under a constant-rate birth-death model. We derive branch length distributions of phylogenetic branches with and without random sampling of individuals of the extant population under two distinct statistical scenarios: a fixed age of the birth-death process and a fixed number of individuals at the time of observation. We find that branches connected to the tree leaves (pendant branches) and branches in the interior of the tree behave very differently under sampling; pendant branches grow longer without limit as the sampling probability is decreased, whereas the interior branch lengths quickly reach an asymptotic distribution that does not depend on the sampling probability.
@@ I Introduction
The reconstructed evolutionary process is central to understanding evolutionary relationships and dynamics. This reconstruction is based on extant genetic data and yields a phylogenetic tree which represents the evolutionary relationships among a set of species, genes, or cells. Crucially, all lineages that eventually went extinct (or were not sampled) are missing from such a tree. This affects the statistics of the reconstructed tree: Under a stochastic birth-death process, species go extinct at a specific rate (or cells die, in a cellular system). Looking at the reconstructed tree, however, only speciation events are seen. To compare empirical phylogenetic trees with models of evolution, it is crucial to understand the link between the underlying evolutionary process and the reconstructed process.
@@
The reconstructed evolutionary process has been studied extensively, with early work of Nee, May, and Harvey Nee et al. (1997) giving the probability density in time for each speciation event in a reconstructed tree from a birth-death process with constant rates. Such a birth-death model describes a population of species, individuals, or cells dividing independently at a fixed rate $\lambda$ and dying independently at a fixed rate $\mu$. For species, birth events correspond to speciations, whereas death events correspond to extinctions. A constant-rate birth-death model serves as a simple model of neutral evolution.
@@
Stadler and collaborators have extended the work by Nee, May, and Harvey Nee et al. (1997) to trees reconstructed from individuals that were sampled from a population Stadler (2009) and to populations conditioned on a particular size Gernhard (2008). The resulting statistics has been used to infer trees and their characteristics Stadler (2013b); MacPherson et al. (2022), and is the cornerstone of widely-used computational packages such as BEAST Bouckaert et al. (2019) and TreeTime Sagulenko et al. (2018). Many extensions of the basic constant-rate birth-death process have been developed, for instance, to account for rates that are variable through time or vary across the tree, see Stadler (2013b) for an overview.
@@
In Stadler and Steel (2012); Mooers et al. (2012) Stadler, Steel and collaborators focus on a particular aspect of tree statistics, namely the distribution of branch lengths in calendar time. They calculate the probability density function for the length of the so-called pendant branches (the lowest branches leading to the leaves of the tree) for trees reconstructed from a given number of leaves. For the particular case of a Yule process (a birth-death process with the rate of death set to zero), they also give the probability density function for the length of interior branches. In Paradis (2016) Paradis introduces an approach to calculate branch length distributions of both pendant and interior branches at a particular time since the start of the birth-death process.
@@
In this paper, we revisit the problem of the distribution of branch lengths with a particular focus on the trees reconstructed from a finite fraction of extant individuals sampled at the time of observation. We distinguish two distinct statistical scenarios: In the first scenario (scenario i)), the time interval between the start of the birth-death process with a single individual and the time of observation (the end of the birth-death process) is known, but the number of extant individuals at the end of the process is a fluctuating random variable. In the second scenario (scenario ii)), one conditions on the number of individuals at the end of the process, but the time interval between the start and end of the process is unknown.
@@
In Section II, we recapitulate the relevant properties of a constant-rate birth-death process. In Sections III-V we use the approach introduced by Paradis Paradis (2016) to systematically write down concrete expressions for branch length distributions for different types of branches under the first scenario: In Section III we focus on pendant branches, in Section IV on branches one and two levels above a pendant branch, and in Section V on generic branches in the interior of a phylogenetic tree. Sections III and V coincide with the work of Paradis in Paradis (2016); they are included here for completeness and because Paradis (2016) does not give concrete expressions for the branch length distributions. Section VI gives the key results of this paper; we consider branch lengths arising in trees reconstructed from only a finite fraction of the extant population. We find that sampling affects different parts of the reconstructed tree differently: while pendant branches become longer and grow without limit as the sampling probability is taken to zero, the distribution of interior branch lengths quickly reaches an asymptotic limit as the sampling probability decreases. In Section VII we turn to the second scenario (conditioning on the number of extant individuals at the time of observation) and introduce a novel approach based on tracking the birth-death process first backwards in time from a given number of individuals at the time of observation to the emergence of a particular branch and then again forwards again to the time of observation. For pendant branches, we recover the results of Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) as expected and generalize these results to interior branches.
@@ II Dynamics of a birth-death process
We consider a standard birth-death process where individuals in a population duplicate with a constant rate of birth $\lambda$ and die with a constant rate of death $\mu$. These individuals might describe individual cells (for instance bacteria or tumour cells) undergoing cell division and cell death, individual multi-cellular organisms in a population under asexual reproduction, or species undergoing speciation and extinction events. Throughout, we will refer to individuals undergoing birth and death events for consistency. The dynamics of the number of individuals $n$ is described by a Markov process, with the probability $p_{n}(t)$ of the population consisting of $n$ individuals at time $t$ following the master equation
@@
It turns out to be convenient for the analysis of branch lengths to consider times relative to the time of observation $T$. We denote the length of the interval between the intermediate time $t$ and the observation time $T$ by $\tau_{e}=T-t$. We further define $\tau=T-\tau_{e}$ as the length of the interval between $\tau_{e}$ and the time of the clade’s origin. Specifically, we denote as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{(T)}_{1}(t=T-\tau_{e})$ the probability that a clade started at time $\tau_{e}+\tau$ in the past, and contains at time $\tau_{e}$ in the past only a single individual that could have extant descendants at the time of observation. Again, setting $\tau_{e}=0$ results in $p_{1}^{(\tau_{e}=0)}(\tau)\equiv p_{1}(\tau)$. For the case of $\tau_{e}>0$ the integrating factor becomes $-\int_{0}^{T-\tau_{e}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]=-\int_{0}^{\tau}\operatorname{\mathrm{d}\!}\tau^{\prime}\left[\mu+\lambda\left(1-2p_{0}(\tau_{e}+\tau^{\prime})\right)\right]=\ln\left(\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\right)$ and we obtain
@@ III Branch length distribution: pendant branches
The results from section II can be used directly to determine the statistics of branch lengths of a phylogenetic tree. In this section, we apply them to the so-called pendant branches, which end in a leaf of the phylogenetic tree. We use the approach of Paradis Paradis (2016) to determine branch length distributions and write down concrete expressions for the probability densities. This section and the following are concerned with complete sampling (where the phylogenetic tree is constructed from the entire population alive at a particular time since the start of the birth-death process), in Section VI we will turn to the case where only a finite fraction of the population is available to reconstruct a tree from.
@@
For a finite time since the start of the birth-death process, $\tau$ has a finite support and hence $P_{\text{pend}}(\tau)$ cuts off to zero when $\tau$ is larger than the time $T$ since the start of the birth-death process. The normalizing factor of (10) is then found to be $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(\lambda(\lambda-(\lambda-2\mu)\operatorname{\mathrm{e}}^{-2(\lambda-\mu)T}-2\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))$. We focus on the particularly simple case of an asymptotically long running time $T\to\infty$. In this case, the cutoff moves to infinity and we find the normalizing factor in (10) to be one, making (10) a proper probability density function. Since the normalizing factor converges exponentially towards one, the asymptotic distribution is also a valid approximation for populations of finite age $T$ as long as $(\lambda-\mu)T\gg 1$. We will focus on this asymptotic case also when treating branches in the tree’s interior.
@@ IV Branch length distribution: interior branches above pendant branches
Equation (11) gives the joint probability density function of branch lengths of pendant and level-two branches up to a normalization factor, which again turns out to be one. Figure 2D compares this result to numerical simulations. A population with birth rate $\lambda=1$ and death rate $\mu=0.75$ was grown from a single individual until time $T=25$ (where the population contained $4460$ individuals), and the phylogenetic tree was reconstructed. The lengths of level-two branches were collected in a histogram and compared to the joint probability $P_{\text{level}}(\tau_{1},\tau_{2})$ given by (11) marginalized with respect to $\tau_{1}$.
@@ V Branch length distribution: interior branches
So far, we have again only considered the asymptotic case of $T\rightarrow\infty$. For finite $T$, the probability of $\tau_{s}>T$ is 00, resulting in a normalization factor of $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(a(1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))^{2}$ for (13). Again, this factor decays exponentially towards one as the population age $T$ increases, justifying the asymptotic results also as an approximation for populations of finite age with $(\lambda-\mu)T\gg 1$.
VI Finite sampling
@@
$P_{\text{int}}^{(\rho)}(\tau)$ $=(\lambda-\mu)^{2}\frac{2\lambda\rho-\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left(\mu-\lambda(1-\rho)\right)}{\left(\lambda(1-\rho)-\mu+\lambda\rho\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}\right)^{2}}\ .$ (20)
@@
ARefer to caption B Refer to caption
Figure 3: Pendant and interior branches at different sampling probabilities $\rho$. (A) We compare the analytical results for the probability density of the length of pendant and interior branches under sampling ((18) and (20) respectively, solid blue lines) to numerical simulations. For the latter, we show the distribution of branches from 100 trees. For every tree, we simulated the growth of a population with $\lambda=1$ and $\mu=0.75$ over time $T=40$, restarting when the population died out or the final population size was not within $2\times 10^{4}$ and $3\times 10^{5}$ (grey bars). The first row of plots shows the branch length distribution of the full trees. For the next row of plots, individuals were sampled with probability $\rho=0.1$ and the corresponding trees were reconstructed. For the third row, each of the previous samples was sampled again with probability $0.1$, corresponding to an overall sampling probability of $\rho=0.01$. The same was repeated for the forth row of plots, yielding $\rho=0.001$. Every tree contributes equally to the empirical distributions depicted. The first column shows the distributions of pendant branch lengths, which shift to longer branches as $\rho$ decreases. The second column of plots shows the distribution of interior branches, quickly reaching the asymptotic distribution (21) (red dashed line) as $\rho$ becomes small. (B) A toy example of how sampling affects the length of pendant branches, see text.
@@
The difference between pendant and interior branches under sampling has a simple qualitative explanation, which we explore within a toy problem. Consider a reconstructed tree which is perfectly balanced; both clades defined by each bifurcation have the same size. An example is shown in Fig. 3B. Then a fraction $\rho$ of the extant nodes is sampled. Under sampling, a pair of pendant branches linked to a shared ancestor is defined by the birth of a clade (shown in blue in Fig. 3B) whose size $n$ at the time of observation is defined by $n\rho\approx 2$ (so two of its members are expected to be sampled). The number of branch levels between this founding birth event and the time of observation is $\ln_{2}(n)$, so the pendant branch lengths are increased under sampling by a factor of $1-\ln_{2}\rho\approx-\ln_{2}\rho$ relative to the fully sampled case. (This argument assumes that on the fully sampled reconstructed tree, pendant and interior branches are approximately of the same length.) On the other hand, branches that are interior in the reconstructed tree (i.e. lie above birth event shown in blue in Fig. 3B) have their lengths unchanged by sampling. The distribution of interior branches only changes because under sampling, fewer and older branches are counted as interior. Older branches tend to be a little longer (see equation (13)), because any additional lineages that arose along them had more time to die out until the time of observation.
To show this quantitatively, we compute the cumulative distribution function of the end $\tau_{e}$ of interior branches
@@
$\int_{0}^{\tau_{e}}\operatorname{\mathrm{d}\!}\tau^{\prime}_{e}\int_{\tau^{\prime}_{e}}^{\infty}\operatorname{\mathrm{d}\!}\tau_{s}P_{\text{int}}^{(\rho)}(\tau_{s},\tau^{\prime}_{e})$ $=\lambda\rho\frac{1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}}}{\lambda\rho+\left(\lambda(1-\rho)-\mu\right)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}}}\ .$ (22)
For any finite $\tau_{e}$, this distribution approaches zero with decreasing sampling probability $\rho$. This means that the ends of interior branches lie asymptotically far away from the time of observation when $\rho$ approaches zero. Setting equation (22) equal to $1/2$ and solving for $\tau_{e}$ yields the median interior branch end. In the regime $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}}\ll 1$, the median is given by $\ln\left((\lambda-\mu)/\rho\right)/(\lambda-\mu)$, reproducing the dependence on the sampling probability found in the toy model.
VII Branch lengths conditioned on the population size at the time of observation